Consider the following model of the labor market: Labor supply: [ = āxw+7 Labor demand: 24 = f - w The values of the equilibConsider the following model of the labor market:
Answer #1

habor supply: L = axw té ; 2d = fw At equilibrium (d = L = axho tł = gw ăw tê= 5- w anw tw/f-ł äwtw if- whax+12= f - ł - wlat